• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name quemener

    Log entry time 16:09:24 on March18,2000

    Entry number 36456

    keyword=shift summary

    Shift Summary:

    08h00 Access in order to repair the hadron arm VDC. Bogdan changed the amplifier card on the hadron arm bottom VDC. Taking two cosmics runs at the same time.
    Bogdan replaced a TDC.
    10h00 Michael Kuss called and mentionned that he has seen no EPICS data
    related to the BPM after 2/3 of run 2805 and in run 2806. Bob Michaels is also looking into that problem. The reboot of CODA after the TDC change should have taken care of that.
    During the access, Bogdan increased a little bit the high voltages on Hadrom arm for:
    S2/ S3 L + 125 V
    S2/ S4 R + 50 V
    S2/ S5 R + 35 V
    10h35 We asked for beam and we did Harp scans:
    11h09 OTR measurement at 5 uA: 0.310 mm for x_width and 0.106 mm for y_width.
    at 30 uA 0.294 mm for x_width and 0.112 mm for y_width. At this energy, this is acceptable.
    11h36 Mcc called, some problems in the main machine under investigation.
    12h00 beam is back we are finally going to take data...
    15h30 reboot of the target, last minutes of run 2818 should be removed (about 10000 events)